# Fatigue-PINN: Physics-Informed Fatigue-Driven Motion Modulation and Synthesis

### v1

@@ IV-A Fatigue Effects in Animation
IV-B Accuracy Results
@@
In Table I we provide the Normalised Root Mean Squared Error (NRMSE) values for joint-specific ID and FD models of Fatigue-PINN framework across different actions. NRMSE values indicate the discrepancies between ground truth and predicted joint angles/torques during the testing of the models, thus, smaller values signify good accuracy. As observed in Table I all values are $<1$ , consequently, our models perform quite well in new (unseen) data.
TABLE I: NRMSE For ID/FD Models
@@
Punching $0.17$/$0.72$ $0.11$/$0.90$ $0.91$/$0.36$ $0.75$/$0.13$ $0.16$/$0.37$ $0.27$/$0.99$
Throwing $0.49$/$0.58$ $0.12$/$0.57$ $0.28$/$0.69$ $0.39$/$0.62$ $0.14$/$0.56$ $0.19$/$0.74$
Waving $0.21$/$0.24$ $0.23$/$0.24$ $0.11$/$0.72$ $0.31$/$0.25$ $0.93$/$0.19$ $0.58$/$0.14$
@@
Punching $0.19$/$0.20$ $0.23$/$0.79$ $0.37$/$0.55$ $0.17$/$0.28$ $0.45$/$0.71$ $0.57$/$0.63$
Throwing $0.10$/$0.57$ $0.20$/$0.62$ $0.39$/$0.65$ $0.73$/$0.50$ $0.80$/.79.79.79.79 $0.13$/$0.67$
Waving $0.35$/$0.32$ $0.41$/$0.15$ $0.10$/$0.18$ $0.40$/$0.20$ $0.27$/$0.12$ $0.34$/$0.25$
@@
Punching $0.22$/$0.15$ $0.29$/$0.28$ $0.35$/$0.30$ $0.16$/$0.31$ $0.25$/$0.24$ $0.16$/$0.32$
Throwing $0.99$/$0.12$ $0.19$/$0.57$ $0.71$/$0.44$ $0.38$/$0.58$ $0.66$/$0.71$ $0.52$/$0.58$
Waving $0.37$/$0.20$ $0.36$/$0.30$ $0.19$/$0.15$ $0.20$/$0.29$ $0.30$/$0.39$ $0.59$/$0.64$
  • a
    NRMSE values for joint torques/angles of the right upper and lower limb as well as torso, across punching, throwing, and waving motions. The first values in each row indicate the NRMSE values produced by the joint-specific ID models while the second values (after the ”/”) denote the NRMSE values for joint-specific FD models.
To further support the selection of a joint-specific approach, i.e. apply Fatigue-PINN for each joint angle, against the non-joint-specific, we provide the Pearson’s Correlation Coefficient, $R^{2}$, values to evaluate the precision of the two methods. $R^{2}$ evaluates the correlation between the predicted joint torques/angles and their real values, with values closer to $1$ indicating better alignment. As shown from Table II the predictions of the ID and FD models of the joint-specific architecture tend to be more correlated (higher $R^{2}$) with the ground-truth values, than the ones of the non-joint-specific approach, providing an insight of average testing accuracy for the former method being $\sim 97\%$, in contrast of average testing precision for the non-joint-specific, $\sim 86\%$. As expected, the joint-specific Fatigue-PINN architecture enhances the precision of animations by leveraging neural networks to extract latent space correlations from all joint torques/angles in order to efficiently accomplish a single task. It is worth noting that the $R^{2}$ values are computed for the joint torques/angles of the upper right limb and during punching motion. Similar results with minimum differences $\sim 1-2\%$ are observed in the other two movements as well as the rest of the joints.
TABLE II: $R^{2}$ For Joint-Specific and Non-Specific Approaches
@@
Motion arm_flex arm_add arm_rot elbow_flex elbow_pro wrist_flex
JS Model $0.97$/$0.96$ $0.90$/$0.91$ $0.96$/$0.95$ $0.94$/$0.96$ $0.93$/$0.97$ $0.94$/$0.95$
NJS Model $0.88$/$0.90$ $0.81$/$0.85$ $0.84$/$0.88$ $0.85$/$0.87$ $0.82$/$0.89$ $0.81$/$0.82$
  • b
    Pearson’s Correlation Coefficient values ($R^{2}$) for predictions of ID (estimating joint torques) and FD (estimating joint angles) models, during punching motion. A comparison of $R^{2}$ values between the Fatigue-PINN joint-specific approach and the non-joint-specific approach is provided. Thus, JS stands for ”Joint-Specific”, and NJS, for ”Non-Joint-Specific” models. Just as in I, $R^{2}$ values of the ID models are presented first, while the second values (after the ”/”) denote the $R^{2}$ values for the respective FD models.

### v2

@@ IV-A Fatigue Effects in Animation
IV-B Perceptual Study
To further evaluate the fatigued motion sequences produced by Fatigue-PINN in terms of conveying externally perceived fatigue, we conducted a perceptual study. The study included a questionnaire drafted based on the existing literature [72, 73, 74] that describes good practices for performing visual quality evaluations on animation videos, and especially on the work of the Disney Research Team [75], where edited throwing animations were assessed in terms of perceptual plausibility. In accordance with the above works, we constructed a questionnaire in which participants were exposed to short videos (of duration $\sim 30-40$ seconds), showing fatigued (depicting various levels of fatigue) and non-fatigued movements, and were asked to:
  • identify which of these videos they perceive as the most fatigued,
  • specify which fatigued effects they consider are conveyed through the fatigued animation videos,
  • assess the fatigued motion sequences in terms of realism/physical-plausibility, as well as
  • provide their feedback regarding what they expect to see from a fatigued motion of X action (i.e. punching, throwing, or waving).
Our study group consists of $19$ participants, among which $13$ ($68.4\%$) are male and 66 ($31.6\%$) are female, while a ”Prefer not to say” option was included to preserve the right of the participant not to identify with a binary gender. The age group of the participants is 184418-44, with more than half ($11$ - $57.9\%$) of the participants belonging in the range of 253425-34 years old. We also asked the participants to specify their experience in research areas relevant to this work, namely Animation, Computer Graphics, and Biomechanics, to obtain feedback from both inexperienced and individuals who are knowledgeable in motion synthesis tasks. Hence, $10$ ($52.6\%$) out of $19$ participants specified to have experience in the Computer Graphics area with $5$ of them being experienced in both Animation and Computer Graphics, 22 ($10.5\%$) are working in the Biomechanics area, 11 ($5.3\%$) participant is experienced only in Animation, and 66 ($31.6\%$) are inexperienced. Out of $13$ participants being experienced in the above research areas, 11 ($5.3\%$) individual claimed to have ”Excellent” knowledge in their area, 66 ($31.6\%$) claimed to have ”Good” experience, and the remaining 77 ($36.8\%$) have an ”Average” level of experience in their area. The level of expertise was specified on a scale from ”Excellent” to ”Average”, while individuals who do not have any experience in the mentioned research areas selected the ”None” option.
In table II, the percentage of participants selecting the fatigued motion sequence as the one they perceive as the most fatigued versus non-fatigued or less fatigued movements, is illustrated:
TABLE II: Percentage of Participants Correctly Perceiving the Most Fatigued Motion
Motion Animation Video Trial Percentage $\chi^{2}$ P-value
Punching NF VS $30\%$ F $94.7\%$ 0.00010.0001*
Waving $30\%$ F VS $10\%$ F $84.2\%$ 0.00290.0029*
Throwing $20\%$ F VS $40\%$ F $73.7\%$ 0.03890.0389*
Waving $20\%$ F VS $40\%$ F $89.5\%$ 0.00060.0006*
Punching NF VS $20\%$ F $94.7\%$ 0.00010.0001*
Punching $50\%$ F VS $30\%$ F $89.5\%$ 0.00060.0006*
Throwing $10\%$ F VS $50\%$ F $84.2\%$ 0.00290.0029*
Punching $40\%$ F VS NF $89.5\%$ 0.00060.0006*
Throwing $0\%$ F VS $30\%$ F $78.9\%$ 0.01160.0116*
Punching NF VS $40\%$ F VS $20\%$ F $73.7\%$ 0.03890.0389*
Punching $10\%$ F VS $20\%$ F VS $30\%$ F $63.2\%$ $0.1083$
Throwing $40\%$ F VS $20\%$ F VS $0\%$ F $47.4\%$ $0.4737$
Punching $0\%$ F VS $40\%$ F VS $20\%$ F $73.7\%$ 0.03890.0389*
Waving $20\%$ F VS $0\%$ F VS $40\%$ F $73.7\%$ 0.03890.0389*
Throwing $50\%$ F VS $0\%$ F VS $10\%$ F $57.9\%$ $0.2513$
  • b
    Various fatigued motion sequence comparisons were provided to the study participants in the form of animated videos via the questionnaire. The table illustrates the percentage of responders accurately perceiving the motion with the highest level of fatigue among others of the same motion type (either punching, throwing, or waving). In the last column, the $p$-values derived from the $\chi^{2}$ test are reported. $P$-values smaller than significance level $\alpha=0.05$, are marked with an asterisk (*). ”NF” and ”F” stand for ”Non-Fatigued and ”Fatigued” motion, respectively.
As illustrated in Table II (see column ”Animation Video Trial”), we opted to create video sequences that depict both huge discrepancies in motion — such as a comparison between $10\%$ and $50\%$ fatigue-induced movements —as well as more subtle variations (e.g. motion sequences being $10\%$ fatigued versus $20\%$ fatigue). Furthermore, we randomized the order of the non/less/more fatigued motions in each animation trial, meaning that the most fatigued motion may appear on the right, left, or middle of the screen, to minimize participant bias. The participants were not provided with any additional information regarding the questionnaire, apart from its main aim; to perceptually evaluate fatigued human motion sequences.
Based on the findings from II (refer to ”Percentage” column), in the majority of cases, participants accurately perceived the animation trial that exhibited the most pronounced fatigue effects. As was expected, in cases where small differences in motion sequences are depicted, the percentage of participants correctly identifying the most fatigued motion is decreasing. For instance, in the trial where punching motion sequences of $10\%$, $20\%$, and $30\%$ of fatigue were compared, $12$ out of 1919 ($63.2\%$) participants detected the motion with the most fatigue effects. Furthermore, the throwing motion seems to have confused the participants the most, with only $9$ participants ($47.4\%$ of the study group) accurately pinpointing the most fatigued motion in an animated video containing motion sequences fatigued up to $20\%$ and $40\%$ as well as a non-fatigued throwing motion. One reason behind this is due to not providing the participants with multiple view angles of the same motion to aid them in detecting differences in the motions as the ones illustrated in Fig. 7 and those analyzed in Section IV-A. Overall, asking responders to compare more than two fatigued motion sequences appeared to be challenging for them.
The participants were also asked to define the fatigue effects that they perceived as being conveyed through the fatigued animated sequences, for each motion type. Notably, this query was deliberately placed at the end of the questionnaire to prevent any potential bias, ensuring that participants did not consciously look for specific fatigue-related elements in the animated fatigued motions. The participants could choose more than one fatigue effect, while the option ”Other” was added to the query to enable participants to add their own observations that made them choose the respective motion as the most fatigued one. Our findings regarding this query are outlined below:
  • Punching Motion: Most of the participants were able to identify fatigue-induced elements within the fatigued motion sequences, such as i) the character’s punches ”missing” their target ($12$ - $63.2\%$ of participants), ii) the character being ”unguarded” to hits ($9$ - $47.4\%$ of participants) and the character’s arm performing a wider movement while hitting ($11$ - $57.9\%$ of participants). These observations are aligned with the results analyzed in Section IV-A. On the contrary, some participants detected elements that are not present in the respective motions, such as the character’s arm performing a more narrow movement while punching ($5$ - $26.3\%$ of participants). The same participants paired the latter observation with falsely assuming that the most fatigued motion sequence was the non-fatigued one in previous questions of the questionnaire.
  • Throwing Motion: As mentioned above, the participants struggled with both specifying the throwing animation trial that conveyed the most fatigued effects and the effects themselves. A small subset of the responders recognized some of the fatigue elements outlined in Section IV-A, including the character throwing an invisible object to a different direction compared to the non-fatigued motion ($4$ - $21.1\%$ of participants), the character being inclined to the non-throwing side ($3$ - $15.8\%$ of participants) as well as the character’s shoulder being positioned higher in the fatigued movement ($4$ - $21.1\%$ of participants). Moreover, the participants provided their own key observations, perceiving the character as being ”out of balance” ($8$ - $42.1\%$ of participants), while unique responses were provided by individual participants ($5.3\%$ of participants) highlighting that the character’s arm appeared ”more tired and lower (placed) as it should” and ”less range of motion, failure to perform the whip-like movement”, was noted. One responder stated that ”all looked exactly the same”, stressing their inability to distinguish between the different levels of fatigue within the throwing motion animated videos.
  • Waving: For the waving motion, the majority of responders ($13$ - $68.4\%$ of the participants) were able to discern that the waving hand exhibits a limited range of motion in the fatigued motion, while $11$ participants ($57.9\%$) perceived that ”the character struggles to perform the task”.
Using a five-point Likert scale - where $1$ corresponds to ”Very Unrealistic” and $5$ to ”Very Realistic” — participants were asked to evaluate the extent to which the fatigued motions they were exposed to convey the externally perceived fatigue. The majority of responders characterized the fatigued-induced motion that our Fatigue-PINN model produces as ”realistic” across all motion types. In particular, $12$ participants ($63.2\%$) found the fatigued punching motion ”realistic”, while the rest of the results were formed as follows: 22 ($10.5\%$) participants found the fatigued punching motion as ”unrealistic”, 44 ($21.1\%$) as ”neutral” and 11 ($5.2\%$) as ”very realistic”.
Nine ($47.4\%$) of the participants described the throwing motion as ”realistic”, while 66 ($31.6\%$) responders characterized it as ”Neutral”, being in line with their responses in previous questions (i.e. not differentiating between fatigued and non-fatigued throwing motions). The realism ratings for the fatigued throwing motion sequences are complemented with 33 ($15.8\%$) participants describing them as ”unrealistic” and 11 ($5.2\%$) as ”very realistic”. No ”very unrealistic” ratings were reported for the fatigued punching and throwing motions. The fatigued waving motion was also described as ”realistic” by more than half of the study group ($10$ participants - $52.6\%$), while 33 ($15.8\%$) participants commented on the motion being ”very realistic”. On the contrary, another 33 ($15.8\%$) responders found the fatigued waving movements ”unrealistic”, 11 ($5.2\%$)”very unrealistic”, and 22 ($10.5\%$) as ”neutral”.
We also requested the feedback of the participants regarding what fatigue elements they expect to see from either a punching, throwing, or waving fatigued motion. In the case of a fatigued punching motion sequence, some of the participants ($8$ - $42.1\%$) commented on fatigue elements already present in our produced animations, such as ”lower positioning” and ”wider movement” of the hands, as well as ”less consistent moves” (i.e. ”missing target”, ”struggling to hold up arms”), while more than the half of the study group ($10$ - $52.6\%$) were expecting to see ”slower motion”. Other interesting remarks consist of ”bad balance, bad posture”, ”mechanical representation of weight (hunched shoulders, bad posture, heaving)” as well as ”legs should also display more fatigue, like less or more restricted movement”. Thus, we conclude that the participants were not able to discern alterations in the character’s posture such as in lumbar bending and rotation rendering the character bending more to the left side while hitting with the right hand, or in the torso’s rotation resulting in turning more to the right, as described in Section IV-A. Furthermore, it was expected for the responders to observe the absence of fatigue in the legs, since in the context of the movements investigated in this work, we opted to apply fatigue only on the upper body.
Regarding the fatigued throwing motion, similar comments to the ones received for the punching one, were collected from the participants, including observations from the motion produced via Fatigue-PINN (”wider motion of the hands”, ”hands being placed lower”, ”leaning”, ”discrepancies in the position and rotation of the shoulder and elbow to fulfill the motion” etc.) and other effects of externally perceived fatigue (e.g. ”bad balance, maybe slipping”, ”limited range of motion”, ”fatigue also in legs”, ”lacking full arm extension”). The presence of fatigue in the character’s torso was also highlighted by the participants, but they anticipated to perceive more intense fatigue effects during throwing movements (e.g. ”I would expect that the character would use its torso more instead of the arm.”). As for the waving motion, the participants reported expecting to see the following fatigue elements in a fatigued waving motion: ”elbow flexion featuring a narrow range of motion”, (similarly) ”waving angle to be shorter”, ”difficulty to wave”, ”a slighter raise of the hand” and ”bad posture”. However, 22 ($10.5\%$) of the participants, were not able to determine such fatigue effects in waving motion ”without context”, and commented that they perceive the animated results depicted in the videos within the questionnaire to be evident in cases where the levels of fatigue are relatively high (”less proclaimed fatigue in general at easier tasks like waving”). In all motion types, $1$ or $2$ participants were expecting to also see different ”facial/sad expressions” to complement the overall sentiment of fatigue, which, however, is out of the scope of this work.
We hypothesize that by providing participants with longer-in-duration animated videos as well as videos of the same motion at different frame speeds (”slow-motion” videos, etc.), the participants may be able to detect more fatigued-induced characteristics in the respective motions, not ”missing” discrepancies in the motion of the torso and shoulders.
As for statistical analyses, we performed a chi-square ($\chi^{2}$) Test of Independence at significance level $\alpha=0.05$, to examine whether there is a significant association between the participants’ selections of the most fatigued motion sequences and the expected results (“actual” fatigued motions). The null hypothesis states that no such relationship exists, and a $p$-value smaller than $\alpha$ indicates the rejection of the null hypothesis. Observing Table II (see column $\chi^{2}$ P-value”) in $12$ out of $15$ tests, $p<0.05$, meaning that the null hypothesis is rejected, hence participants opted for the most fatigued motion sequence accurately (i.e. there is a significant correlation between the participants’ perception of fatigue across fatigue levels and the ”correct” answers). The cases where $p>0.05$ were the ones where participants struggled to correctly discern fatigue differences in the animated motions, resulting in non-statistically significant outcomes.
Finally, a One-Way Analysis of Variance (ANOVA) was utilized to determine whether the different fatigue levels illustrated in the animated videos provided to the participants led to significantly different realism ratings obtained through Likert scales. The null hypothesis in this case states that no statistically significant differences in realism ratings are present, i.e. responders gave consistent realism evaluations, which is a desirable outcome. As can be deduced from Table III, all $p$-values >=0.05>=0.05, in which case the null hypothesis is not rejected, meaning that the realism perception of the study participants does not differ significantly while assessing motion sequences conveying different levels of fatigue. Therefore, the participants consistently characterizing the fatigued motions as ”realistic” is a statistically significant result.
TABLE III: ANOVA $P$-values for Fatigue Level Realism Evaluations
Punching Motion Throwing Motion Waving Motion
pp-value $0.1782$ $0.3630$ $0.4254$
  • c
    The $p$-values as arisen by One-Way ANOVA tests performed to specify whether the realism perception of participants while assessing fatigued motion sequences differs significantly across the fatigue levels of the motions.
Both statistical analyses were implemented using built-in Python functions.
IV-C Accuracy Results
@@
In Table IV we provide the Normalised Root Mean Squared Error (NRMSE) values for joint-specific ID and FD models of the Fatigue-PINN framework across the motion trials of an action. These NRMSE values are average values computed from the results of the three validation folds for each motion type, namely waving, punching, and throwing. NRMSE values indicate the discrepancies between ground-truth and predicted joint angles/torques during the testing of the models, thus, smaller values signify good accuracy. As observed in Table IV all values are <1<1, consequently, our models perform quite well in new (unseen) data.
TABLE IV: NRMSE For ID/FD Models
@@
Punching 0.170.17/$0.72$ 0.110.11/$0.90$ 0.910.91/$0.36$ 0.750.75/$0.13$ 0.160.16/$0.37$ 0.270.27/$0.99$
Throwing 0.490.49/$0.58$ 0.120.12/$0.57$ 0.280.28/$0.69$ 0.390.39/$0.62$ 0.140.14/$0.56$ 0.190.19/$0.74$
Waving 0.210.21/$0.24$ 0.230.23/$0.24$ 0.110.11/$0.72$ 0.310.31/$0.25$ 0.930.93/$0.19$ 0.580.58/$0.14$
@@
Punching 0.190.19/$0.20$ 0.230.23/$0.79$ 0.370.37/$0.55$ 0.170.17/$0.28$ 0.450.45/$0.71$ 0.570.57/$0.63$
Throwing 0.100.10/$0.57$ 0.200.20/$0.62$ 0.390.39/$0.65$ 0.730.73/$0.50$ 0.800.80/.79.79 0.130.13/$0.67$
Waving 0.350.35/$0.32$ 0.410.41/$0.15$ 0.100.10/$0.18$ 0.400.40/$0.20$ 0.270.27/$0.12$ 0.340.34/$0.25$
@@
Punching 0.220.22/$0.15$ 0.290.29/$0.28$ 0.350.35/$0.30$ 0.160.16/$0.31$ 0.250.25/$0.24$ 0.160.16/$0.32$
Throwing 0.990.99/$0.12$ 0.190.19/$0.57$ 0.710.71/$0.44$ 0.380.38/$0.58$ 0.660.66/$0.71$ 0.520.52/$0.58$
Waving 0.370.37/$0.20$ 0.360.36/$0.30$ 0.190.19/$0.15$ 0.200.20/$0.29$ 0.300.30/$0.39$ 0.590.59/$0.64$
  • d
    NRMSE values for joint torques/angles of the right upper and lower limb as well as torso, for punching, throwing, and waving motions. These NRMSE values are derived by averaging the prediction results across all validation folds of each action. The first values in each row indicate the NRMSE values produced by the joint-specific ID models, while the second values (after the ”/”) denote the NRMSE values for joint-specific FD models.
To further support the selection of a joint-specific approach, i.e. apply Fatigue-PINN for each joint angle, against the non-joint-specific, we provide the Pearson’s Correlation Coefficient, $R^{2}$, values to evaluate the precision of the two methods. $R^{2}$ evaluates the correlation between the predicted joint torques/angles and their real values, with values closer to $1$ indicating better alignment. As shown from Table V the predictions of the ID and FD models of the joint-specific architecture tend to be more correlated (higher $R^{2}$) with the ground-truth values, than the ones of the non-joint-specific approach, providing an insight into average testing accuracy for the former method being $\sim 97\%$, in contrast of average testing precision for the non-joint-specific, $\sim 86\%$. As expected, the joint-specific Fatigue-PINN architecture enhances the precision of animations by leveraging neural networks to extract latent space correlations from all joint torques/angles in order to efficiently accomplish a single task. It is worth noting that the $R^{2}$ values are computed for the joint torques/angles of the upper right limb and during punching motion. Akin to the computation of the NRMSE values, the $R^{2}$ values are also average values calculated from the predictions of the three validation folds of an action (in our case the punching motion). Similar results with minimum differences $\sim 1-2\%$ are observed in the other two movements, as well as the rest of the joints.
TABLE V: $R^{2}$ For Joint-Specific and Non-Specific Approaches
@@
Model arm_flex arm_add arm_rot elbow_flex elbow_pro wrist_flex
JS Model 0.970.97/$0.96$ 0.900.90/$0.91$ 0.960.96/$0.95$ 0.940.94/$0.96$ 0.930.93/$0.97$ 0.940.94/$0.95$
NJS Model 0.880.88/$0.90$ 0.810.81/$0.85$ 0.840.84/$0.88$ 0.850.85/$0.87$ 0.820.82/$0.89$ 0.810.81/$0.82$
  • e
    Average (across all validation folds of a motion type) Pearson’s Correlation Coefficient values ($R^{2}$) for predictions of ID (estimating joint torques) and FD (estimating joint angles) models, during punching motion. A comparison of $R^{2}$ values between the Fatigue-PINN joint-specific approach and the non-joint-specific approach is provided. Thus, JS stands for ”Joint-Specific”, and NJS, for ”Non-Joint-Specific” models. Just as in IV, $R^{2}$ values of the ID models are presented first, while the second values (after the ”/”) denote the $R^{2}$ values for the respective FD models.
In order to validate quantitatively the visual results discussed in Section IV-A, we compute the $R^{2}$, and L2 metrics to measure the discrepancies in joint angles before and after applying a certain level of fatigue on each one of the different motion types (punching, throwing and waving). The values of these metrics are reported in Table VI, for the shoulder flexion, shoulder adduction, elbow flexion, lumbar bending, and lumbar rotation joint angles, whose fatigue-induced alterations in motion are mostly discussed in Section IV-A. L2 provides a quantification of the ”distance” between joint angles before and after injecting fatigue in motion. As shown in Table VI, increases in the level of fatigue imply higher L2 values, meaning greater differences in joint angles, and hence more profound visual fatigue effects. The latter observation is valid for all types of motion, while $R^{2}$ values remain relatively high, indicating that there is still a correlation between the non-fatigued and fatigued data. It is worth noting that some angles, like shoulder flexion, are more affected by fatigue compared to others. Concluding, the throwing motion exhibited smaller differences in joint angles across fatigue levels compared to the other two motions, justifying why the study participants were not able to perceive evident changes due to fatigue in throwing motion sequences.
[…]
Punching Motion
Motion arm_flex arm_add elbow_flex lumbar_bend lumbar_rot
NF VS $20\%$ F 2.012.01/$0.92$ 2.712.71/$0.90$ 1.981.98/$0.92$ 1.361.36/$0.94$ 2.632.63/$0.90$
NF VS $30\%$ F 3.213.21/$0.90$ 4.244.24/$0.88$ 2.072.07/$0.92$ 2.122.12/$0.90$ 3.333.33/$0.89$
NF VS $40\%$ F 4.374.37/$0.89$ 5.445.44/$0.87$ 2.402.40/$0.90$ 3.073.07/$0.91$ 4.024.02/$0.90$
NF VS $50\%$ F 5.125.12/$0.88$ 6.236.23/$0.85$ 2.592.59/$0.91$ 3.493.49/$0.90$ 4.384.38/$0.89$
Throwing Motion
Motion arm_flex arm_add elbow_flex lumbar_bend lumbar_rot
NF VS $20\%$ F 1.081.08/$0.96$ 1.071.07/$0.94$ 1.621.62/$0.93$ 0.930.93/$0.95$ 1.041.04/$0.94$
NF VS $30\%$ F 1.501.50/$0.95$ 1.581.58/$0.93$ 1.811.81/$0.93$ 1.141.14/$0.94$ 1.201.20/$0.94$
NF VS $40\%$ F 1.951.95/$0.93$ 2.342.34/$0.92$ 2.442.44/$0.92$ 1.421.42/$0.92$ 1.391.39/$0.93$
NF VS $50\%$ F 2.402.40/$0.90$ 2.752.75/$0.91$ 2.492.49/$0.89$ 1.861.86/$0.92$ 1.931.93/$0.90$
Waving Motion
Motion arm_flex arm_add elbow_flex lumbar_bend lumbar_rot
NF VS $20\%$ F 1.291.29/$0.96$ 0.930.93/$0.97$ 1.181.18/$0.94$ 1.301.30/$0.95$ 1.451.45/$0.93$
NF VS $30\%$ F 1.521.52/$0.96$ 1.391.39/$0.96$ 2.022.02/$0.91$ 1.911.91/$0.95$ 2.342.34/$0.92$
NF VS $40\%$ F 1.811.81/$0.96$ 2.162.16/$0.95$ 2.872.87/$0.91$ 2.662.66/$0.94$ 3.263.26/$0.90$
NF VS $50\%$ F 2.162.16/$0.94$ 2.662.66/$0.94$ 3.003.00/$0.91$ 3.233.23/$0.92$ 3.533.53/$0.91$
  • f
[…]